OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 12 to x = 15.

OpenStudy (hwyl):

so evaluate with the given values of x

OpenStudy (hwyl):

try that first

OpenStudy (astrophysics):

\[\text{average rate of change} = \frac{ f(b)-f(a) }{ b-a }\] x= 12= a, x=15=b

OpenStudy (anonymous):

i dont understand

OpenStudy (hwyl):

do you know what a SLOPE is?

OpenStudy (anonymous):

yes

OpenStudy (hwyl):

For f(x) = 0.01(2)x, find the average rate of change from x = 12 to x = 15 For f(x) = 0.01(2)x is the same as saying y = 0.01(2) x now we have two values of x's provided and we need to evaluate them one at a time \(x_1 = 12 \) and \(x_2 = 15 \) \(y_1 = 001(2) x_1 \rightarrow y_1 = 001(2) *12\) \(y_2 = 001(2) x_2 \rightarrow y_2 = 001(2) * 15\) after we evaluate each y's, we use the SLOPE formula \( \dfrac{y_2-y_1}{x_2-x_1} \) because AVERAGE rate is just the same as the SLOPE.

OpenStudy (hwyl):

there's a period in there 0.01 but it didn't go through my bad.

OpenStudy (anonymous):

so would x1 and x2 and y1 and y2 be the same?

OpenStudy (hwyl):

yes something similar to when you are learning slopes

OpenStudy (anonymous):

so does that mean y1=.24 and y2=.3?

OpenStudy (hwyl):

I do not have a calculator so I am not sure.